Prev Next

SW&IT Meeting Minutes 2020-02-24

Attendees: Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200210
LR20200210 - Bec, Matthieu 80.0
LR20200210 - Holmberg, Carl 72.0
LR20200210 - Edmeades, David 80.0
LR20200210 - Sola, Igor 80.0
LR20200210 - Kubanek, Petr 28.0
LR20200210 - Hooper, Stephen 80.0
LR20200210 - Zhang, Yang 80.0
LR20200210 - Zhang, Xianyu 88.0

News from management

Previous Action Items

Owner Action Due Date Completion Date
DE timecapsule obsdev 2020-02-12  
DE oVirt for PEPSI 2020-02-12  
IS Luci FITS keyword seeing @zenith 2020-02-24  

Open topics

  • PK farewell lunch, AZ room Thu 27
  • Return to fifth floor, Mar 5 ?

Team reports

PK

  • IT cleanup
  • UMAC training
  • web-dms - Docker, Duration option

YZ

  • Updated LBC tech image moving script (and created github pull request to merge).
  • Discussed with Olga for MODS data machine cleaning process.
  • Started to work on a MODS data machine cleaning and data file monitoring script (to be run on MODS data machine, I asked Stephen to mount the LBC repository directories on the MODS data machines).
  • Schedule a time with Olga to test various MODS programs on a 64-bit obs machine (perhaps the "robs" machine).
  • TMS software meeting on Friday.
  • UMAC training
  • Discuss with Petr on TCS issues and troubleshooting.
  • AO-Cravat work restart?
  • Safety training.

DE

  • oVirt ongoing
  • RS1 reboot seems to have solved the twitching
  • Drive replacement march going steadily
  • Parts for summit laptop should all be here today
  • Ordered Mac Mini for obsdev
  • Added hyperlink boxes to IT ticketing page; not ideal but probably the best we'll get with our constraints
  • Archive permissions (ACL) investigation
  • Veeam license install/proxy setup
  • Too many meetings last week to investigate HDMI audio disable on projector; on the list
  • Gentle reminder to Olga re: laptop OS swap
  • Apple has made a new policy for Safari re: security certificates. It will no longer accept any cert generated after 1 Sep that is valid for more than 13 months. Link

MB

  • Last week
    • OC, WPC, Management
    • TMS #2
    • ROWG kickoff meeting (plan for better time, next time)
      • learnt from side discussions about our Iraf use, more as a scripting tool ...
    • Shark-Vis risks
    • DMS m1 / mongodb sub-sampling
    • Finalizing SW post with GG into new system
  • This week
    • ROWG #1 notes cleanup, schedule #2
      • TBC consensus
        • polycom > zoom ?
        • 4|6x heads > 2x 4K ?
    • TMS n/a and planning with Yang
    • TCS ITs review
    • DMS development (push websocket)

XZ

  • Last week
    • DX AdSec log and troubleshotting
    • DX AdSec dicusion and meetings
    • Varsmonitor telemetry recording
  • This week
    • DX AdSec troubleshoting and tests on the mountain

SH

  • Leroy ASC
  • rm507 stability issues
  • KVM into machine room Tucson
  • weather underground
  • ROWG
  • info add DIMM, changes to layout
  • user cleanup in Google and Windows AD
  • displays (split from info, redirection of bandwidth).

next:

  • rm507 stability
  • ovirt
  • ROWG?
  • log copy down

IS

Last week:
  • PIT
    • Release work fixes, in particular resubmission backend
  • LUCI
    • Understanding how mount XML elements affect binocular/synch behaviour
  • Queue Time Accounting
    • fixing bug about consecutive state updates
    • Adding reconfiguration bin

This week:
  • OT
    • Final release?
  • PIT
    • PIT enhancements and release

CH

Last week
  • AHS: Initial draft LBTI alert script, need to add ability to reconnect ot INDI service after dropouts.
  • Support scripts: Improved error handling greatly improved performance of RB_Science. MODS2 still a bit flakier than LBC and MODS1.
  • TCS: Collected AGW3 & 4 following error data while in mountain lab, post-lubrication. Tested AGW1 movement after LUCI1 reinstall.
This week
  • AHS: Add INDI service reconnect and AHS data connection to LBTI alert script.
  • ARGOS: Investigate creating remote GUI client package.
  • Support scripts: Minor investigation into why MODS2 still a bit flakier than LBC and MODS1.
  • TCS: Shadow Petr and Yang on TCS discussions.
Topic revision: r14 - 02 Mar 2020, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback